OpenStudy (anonymous):

Using the equation, y = 20,000(0.95)x, predict the purchasing power of $20,000 ten years later Anwsers: $14,560 $10,255 $11,975 $12,635

OpenStudy (campbell_st):

again this is exponential growth and decay.... y = 20000(0.95)^10

OpenStudy (anonymous):

Umm.. You mean \[20000(0.95)^x\] right??

OpenStudy (anonymous):

again.. this is very difficult for me to figure out..

OpenStudy (anonymous):

$11975

OpenStudy (anonymous):

this stuff is so dumn too me . -_-

OpenStudy (anonymous):

Just plug in 10 for x and put in your calculator

OpenStudy (anonymous):

The equation is y=20,000(0.95)^x So the buying power when x=0 is 20,000(0.95)^0=20,000 The purchasing power of $20,000 ten years later is 20,000(0.95)^10=11975